a car traveling 50km/h east accelerates at 0.25 m/s2 west. What will be its velocity and displacement:

a) at 3min?

b) when it passes it's the initial position?
Answer #1

a) initial velocity is 50 km/h = 13.89 m/s(positive means east)

a = -0.25 m/s^2(negative means towards west)

The displacement when t = 3 min = 180 s is

d= ut + -at = 13.89 x 180 -0.25 x 180/2 = -1550. m = -1.55km

displacement is negative because it is in the west. The magnitude of displacement is d = 1.55 km = 1550 m

velocity after t = 180 s is

v = u + at = 13.89 -0.25 x 180 = -31.11 m/s

In km/h the answer is v = - 112 km/h

velocity is negative because it is in the west direction.

b) When it passes initial position d = 0 m 0 km

Therefore

=

v=-u= -13.89 m/s = -50 km/h

velocity is negative because it is in the west direction.